Sky is completely clouded over, doesn't look to be getting better any time soon.

Pickup noise in images from yesterday seems to have gone away.

3:25pm Start taking darks, starting with 0001

5:11pm Since it's still cloudy, they are going to do some diagnostics on LBC in the meantime

7:09pm Darks finished (through file 0070)

7:22pm Started taking darks for Greg's program starting with 0071, weather still bad, wind speed ~15 m/s with gusts up to 25 m/s

10:09pm There's a chance things could clear up around midnight, but winds are still high (~20 m/s)

11:24pm Greg's darks are done. Going to take some dome flats for imaging, starting with 0136 through 0149

12:02am Still cloudy and windy, not doing anything until conditions improve

2:55am Winds consistently between 15-20 m/s, some holes in the clouds but still looking bad

3:56 AM. We're still closed due to wind and clouds.

-- XiaohuiFan - 03 Mar 2013
Topic revision: r7 - 04 Mar 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback